# On the origins of counterrotating stellar disks in TNG50. I

### v1

@@ 3 Identification of counterrotating disks
Fig. 1 shows the distribution of the counterrotating disk mass fraction obtained from our full sample of late-type disks. Note that, for the majority of disk galaxies, the CRD components do not surpass the $1\%$ threshold, highlighted by the vertical dashed line. We find that only $\approx 8\%$ of the galaxies host a CRD with a mass fraction larger than $1\%$. The sample of galaxies that host a significant CRD, i.e, $M_{\rm CRD}/M_{\rm Disk}\geq 0.01$ consists of 20 galaxies. In this study, we decided to include in our final sample 6 boundary cases with $0.0095\leq M_{\rm CRD}/M_{\rm Disk}<0.01$.
In Table 1 we list the internal properties of the galaxy sample with identified CRDs. In the left most column the label given to each identified CRD is shown. The corresponding Subfind ID at $z=0$ is given in the second column.
Fig. 2 presents 8 examples of galaxies hosting a significant CRD. The figure shows their total stellar surface mass density both, in face-on and edge-on projections. We show the most extensive CRDs in terms of their $R_{95}^{\rm CRD}$. The dashed white circle denotes the overall disk size in terms of $R_{\rm opt}$, while the solid and dot-dashed yellow circles represent $R_{95}^{\rm CRD}$ and $R_{50}^{\rm CRD}$, respectively. We note that the morphological diversity of the galactic disks, with some exhibiting well-defined bars (e.g., CR-5) and spiral structures (e.g., CR-10 and CR-18).
Fig. 3 shows the spatial distribution of the counterrotating stellar particles considered as members of the CRD, in the same order as in Fig. 2. By comparing the two figures, we observe that in some galaxies (e.g., CR-5 and CR-3), the CRDs extend to radial distances comparable to those of the overall stellar disk. In contrast, other cases (e.g., CR-25) exhibit much more compact CRDs. In addition, some cases such as CR-0 and CR-5 display an interesting torus like morphology. We will explore the origin of their structure in the following sections. The full set of 26 galaxies analyzed is shown in Fig. 10 and 11.
From this analysis it follows that, based on our criteria, approximately 10% (including the boundary cases) of the galaxies host a significant stellar CRD. Our results indicate that CRDs are relatively infrequent. This is consistent with previous observational studies, which report CRDs, specially stellar, as rare systems found in only $\approx 1\%$ of MaNGA galaxies (Bao et al. 2022, Gasymov et al. 2025).
@@ 4 Characterization of CRDs
Fig. 4 shows the relationship between the CRD mass fraction, $M_{\rm CRD}/M_{\rm Disk}$ and different measurements of the disk extension and mass distribution. In particular we focus on $R_{50}^{\rm CRD}/R_{\rm opt}$ (top), $R_{95}^{\rm CRD}/R_{\rm opt}$ (middle) and $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}$ (bottom). Data points are color-coded according to the total stellar mass of the disk.
@@
Figure 4: The CRD mass fraction, $M_{\rm CRD}/M_{\rm Disk}$, as a function of different measurements of the counterrotating stellar disk extension and mass distribution for the sample of 26 significant CRD galaxies. Top & middle panels: the radius containing 50% and 95% of the counterrotating stellar disk mass normalized by the optical radius of the galaxy, respectively. Bottom panel: the stellar concentration of the counterrotating disk. All data points are color-coded by the total stellar disk mass. We see that, while most CRDs are compact with the majority having $R_{50}^{\rm CRD}/R_{\rm opt}<0.25$, there are a few extended cases. Overall, we find that our sample of counter rotating galaxies show a variety of structural properties.
@@
The top panel shows that most CRDs are relatively compact in nature, with approximately 88% having half-mass radii ($R_{50}^{\rm CRD}$) smaller than a quarter of their optical radii ($R_{\rm opt}$). However, the middle panel reveals that around 89% of CRDs have $R_{95}^{\rm CRD}/R_{\rm opt}$ values greater than 0.25, and approximately 19% exceed half the optical radii. This indicates that, while most CRDs are compact in terms of their half-mass radii, many have a reasonably extended radial distribution. The bottom panel of Fig. 4 shows a meassure of the CRDs concentration, which is defined as the ratio $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}$. CRDs with $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}\leq 2$ enclose approximately signifcant of their mass in thin ring-like structure. We identify 3 of these torus-like CRDs, which are CR-0, CR-5, and CR-22. Indeed, this can be appreciated on Fig. 3. CRDs with higher concentration values, such as CR-3 and CR-10, exhibit less compact profiles. The wide spread in the CRDs concentration highlights the structural diversity within the sample. Overall, we find that our sample of counter rotating galaxies shows a variety of structural properties. Three CRDs show a ring like structure, some display an extended CRD, while others have a more compact nature with values of $R_{95}^{\rm CRD}$ below half of the disk’s $R_{\rm opt}$.
@@
We now characterize the origin of the stellar particles that conform each CRD. We recall the reader that stellar particles formed from the gas distribution bound to the potential well of the main host galaxy are defined as in-situ. This includes gas that may be recently stripped from closely interacting satellite galaxies. In Fig. 5 we show the fractions of in-situ and ex-situ stellar particles in each significant CRD. The light blue (bottom) fraction of each bar indicates the contribution from in-situ stellar particles, while the other colored fractions indicate the contribution to the ex-situ from different satellite galaxies. Interestingly, a majority of the CRDs present a dominant in-situ contribution, with 19 out of 26 (73%) cases displaying fractions above 80%. The remaining 7 CRDs present a more significant ex-situ contribution; i.e. $20\%$. We find only one CRD dominated by the ex-situ component (CR-6). In general, CRDs with significant ex-situ contributions have only one significant progenitor that contributes over 90% of the ex-situ particles, with CR-2 being the only exception. We will further explore these cases in Section 5.
@@
Figure 5: The fraction of the in-situ and ex-situ stellar particles in the counterrotating disk for the sample of 26 significant CRD galaxies. Light blue bars represent the fraction of in-situ stellar particles in the CRD, while the remaining colors represent the contribution to the ex-situ fraction in the CRD from different satellite galaxies. Each color refers to a different satellite. The red dashed line indicates an in-situ contribution equals to 80%.
@@
In general, we find that CRDs have a very significant in-situ contribution and, in most cases, present a compact distribution. We note however, that some show extended configurations. As we will show in Section 5, even though most CRDs stellar particles are born within the host potential well, the corresponding star formation bursts are typically associated to significant interaction with massive nearby satellites.
@@
In Fig. 6 we show the distribution of CRDs in a space defined by the fraction of ex-situ material and the extension of the CRD with respect to the overall stellar disk, $R_{95}^{\rm CRD}/R_{\rm opt}$. Each symbol represents a single CRD. The top histogram in the figure highlights that a significant fraction of CRDs ($\approx 80\%$) fall into a compact configuration, with values of $R_{95}^{\rm CRD}/R_{\rm opt}<0.5$. Indeed, the median value of the distribution is $R_{95}^{\rm CRD}/R_{\rm opt}\sim 0.37$. Note however the presence of the five interesting cases with very extended in-situ discs; i.e. $R_{95}^{\rm CRD}/R_{\rm opt}>0.6$ (see also Fig. 4). As previously discussed, the histogram on the right panel of the figure shows that most CRDs ($\approx 73\%$) have an ex-situ fraction <0.2<0.2. Indeed, the median value of the ex-situ mass fraction is $\sim 0.06$.
@@
Based on this space we classify CRDs in different categories using a threshold $R_{95}^{\rm CRD}/R_{\rm opt}=0.5$ and an ex-situ fraction of 20%. The different categories are defined as i) compact in-situ, ii) compact ex-situ and iii) extended in-situ. These categories reflect both their spatial extent and whether the ex-situ mass fraction is significant. The classification scheme is highlighted with different colored regions in Fig. 6. Using this scheme, we find that compact in-situ CRDs are the most common, with 14 cases, followed by 7 compact ex-situ and 5 extended in-situ CRDs. Note that no CRDs with a significant ex-situ fraction are found to be extended.
@@
This analysis indicates that CRDs are not only rare in TNG50, but they also predominantly exhibit compact spatial distributions. Interestingly, despite being less frequent, it is the extended types that are more likely to be detected observationally, providing context for the $\approx 1\%$ occurrence reported in the MaNGA survey (Bao et al. 2022, Gasymov et al. 2025). Using this CRD classification as foundation, in the next Section we proceed to analyze each type individually, examining their star formation histories and merger activity to characterize the nature of their formation.
@@
Figure 6: Classification of the sample of 26 significant CRDs according to their counterrotating stellar disk extension ($R_{95}^{\rm CRD}/R_{\rm opt}$), and to their ex-situ fraction of stellar particles. We classify the CRDs into four different types, namely extended ex-situ CRDs (yellow quadrant), compact ex-situ (red rombus quadrant), compact in-situ (blue triangle quadrant) and extended in-situ (orange circle quadrant).
@@ 5 Formation channels for CRDs
In this Section we explore the different formation channels of the 26 significant CRDs, subdivided according the the three categories defined in Section 4. In our analysis we incorporate information about the stellar ages as well as the properties of infalling satellites.
@@
Fig. 7 shows the star formation histories (SFHs) of all 26 significant CRDs. Each panel represents a galaxy from our sample. Here, the orange and the light blue lines represents the SFH of the CR ($\epsilon<-0.7$) and the co-rotating ($\epsilon>0.7$) disk component, respectively. The vertical dashed lines indicate the median of each distribution. Additionally, the gray dashed line represents the SFH of all stellar particles confined within a cylinder of radius $R=R_{\rm opt}$ and height $|z|\leq 5$ kpc. All SFHs are normalized to the peak value of the corresponding distribution. We can clearly see that, with respect to the full disks, CRDs typically have a more bursty history. In most cases, these bursts are also noticeable in the overall disk’s SFH. Additionally, we also notice that CRDs are significantly older than the corotating counterparts. Indeed, this is the case for $\approx 93\%$ of the identified CRDs. The two exceptions are CR-0 and CR-11. We will further study these cases later in this section.
@@
Fig. 7 also highlights the more significant interaction events each galaxy has undergone. We focus on interactions associated with satellites that have at least a total peak mass of $5\times 10^{9}$ M and that, at the corresponding peak mass time, the satellite to host mass ratio was larger than $20\%$. The first pericenter passage of such interactions are indicated with arrows, color coded according to each satellites’s peak mass. For CRDs with an ex-situ fraction $\gtrsim 10\%$, we highlight the arrow associated with the most significant ex-situ contributor. As expected, we observe in all cases a higher frequency of significant mergers at early times. Nonetheless, significant interactions can take place at any time. Such interactions are commonly associated with star-formation (SF) bursts in the overall stellar disk (see also Gargiulo et al. 2017).
5.1 Compact in-situ CRDs
@@
Compact CRDs are the most frequent type in our sample. By definition, these systems are not very extended (their values of $R^{\rm CRD}_{95}$ lying under half of the optical radius of the overall disk), meaning the counterrotating stars are largely confined to the spatial limits of the galaxy’s bulge or bar. Within this category, we first consider those CRDs that formed predominantly in-situ (ex-situ fraction ¡ 20%). There are 14 such cases in our sample. Compact in-situ CRDs are indicated by blue boxes, as well as an upward-pointing blue triangle symbols on the top right corner of each panel in Fig. 7.
@@
Figure 7: The star formation histories (SFHs) of all 26 significant CRDs, grouped according to the classification type, namely compact in-situ, compact ex-situ, extended in-situ (see Sec.4). Here, the orange and the light blue lines represent the SFH of the counterrotating and the co-rotating disk component, respectively. The vertical dashed lines indicate the median of each distribution. The gray line represents the SFH of all stellar particles belonging to the disk. All SFHs are normalized to the peak value of the corresponding distribution. For each CRD, we also label the time of the first pericenter for the most significant interactions experienced by the galaxy, considering only satellites with a peak total mass $5\times 10^{9}M_{\odot}$ and a satellite-to-host mass-ratio ¿20%. The most significant interaction of the CRD is identified by an arrow outlined in black and white.
@@
Among the compact in-situ CRDs, one galaxy, CR-14, stands out as a borderline case: about 18% of its CRD stellar particles are ex-situ (accreted). CR-14 lies just below our threshold for the “ex-situ” category; its counterrotating component can be mainly associated to a significant accretion event. Indeed, Fig. 7 shows that a merger with a satellite of peak mass of the order of $\approx 8\times 10^{10}$ M induces a SF burst that gives rise to most of the co-rotating component. Furthermore, the assembly information shown in Fig. 5 for CR-14’s CRD shows that the majority of its few accreted stars came from this single infalling satellite (see Fig. 7, blue highlighted arrow), which contributed both stars and gas, as well as inducing a burst of star formation.
@@
Many of the compact in-situ CRDs show clear signs of interaction-triggered star formation in their histories, even though the stars themselves formed in-situ. For example, CR-11, CR-12, CR-14, CR-18, CR-20 and CR-25 all exhibit enhanced star formation in their CRDs that was preceded by an interaction event either at intermediate (e.g CR-12 and CR-20) or at late times (CR-11). In these galaxies, a passing or merging satellite disturbed the gas in the host, inducing a burst of star formation in the inner galactic region contributing with stellar particles on retrograde orbits. In each of these cases the CRD’s stellar mass is almost entirely in-situ; the interactions served as a catalyst for star formation rather than directly depositing a large number of stars.
@@
On the other hand, some compact in-situ CRDs appear to have formed very early, alongside the initial assembly of the galaxy’s bulge, and without any later significant interactions. Galaxies such as CR-1, CR-8, CR-9, CR-21, CR-22, and CR-24 exhibit strong early bursts of star formation, occasionally correlated with the high merger rates characteristic of that epoch. During this early phase of galaxy assembly, star formation occurred in a more irregular and turbulent environment, resulting in a pressure-supported stellar population. The more isotropic kinematics of these stellar particles are associated with broad circularity distributions, contributing to both the co-rotating and counterrotating components of the final stellar disk. Accordingly, the stellar ages of these CRDs trace back to the formation of the galaxies’ primordial inner regions. These structures are therefore old and centrally concentrated.
@@
The two remaining cases have peculiar formation mechanisms. CR-16 formed over a prolonged period during which the co-rotating component experienced minimal star formation; until the merger of an infalling satellite, the galaxy lacked a dynamically cold stellar disk. The CRD is thus a remnant of this earlier evolutionary stage, when the galaxy was pressure-supported rather than rotation-supported. Finally, CR-13’s formed within a short time window following the smooth accretion of misaligned cold gas. The accreted gas triggered a burst of SF in the central region giving rise to the CRD.
@@
In summary, the compact in-situ CRDs encompass a variety of formation paths. Some (e.g., CR-1, CR-8, CR-9, CR-21, CR-22, CR-24) formed at the earliest times of their host galaxy, essentially as part of the old bulge formation. Others (e.g., CR-11, CR-12, CR-18, CR-20) formed later, with their counterrotating stars born during one or more bursts triggered by satellite interactions (but without significant stellar accretion). A particular case (e.g., CR-14) shows a minor contribution from accreted stars in addition to in-situ star formation. CR-16’s CRD is the remnant of a pre-disk stellar population formed prior to the development of a dynamically cold disk, while CR-13’s CRD originated from the accretion of misaligned gas that fueled central star formation.
@@
From these results, we can broadly categorize the origins of compact in-situ CRDs into three scenarios: (i) formation at early times together with the galaxy’s bulge, (ii) interaction-induced star formation without substantial stellar accretion, (iii) peculiar cases associated with smooth misaligned cold gas accretion and pre-stellar disk irregular star formation. The first two mechanisms being the most frequent. In practice, some CRDs may involve a combination of these mechanisms. Overall, however, the common theme is that the compact in-situ CRDs were built by star formation within the galaxy, often triggered or influenced by interactions, rather than by direct deposition of stars from mergers.
5.2 Compact ex-situ CRDs

### v2

@@ 3 Identification of coplanar counterrotating disk components
Fig. 2 shows the distribution of the CRD mass fraction obtained from our full sample of late-type disk galaxies. Note that for the majority of disk galaxies, the CRD components do not surpass the $1\%$ threshold, which is highlighted by the vertical dashed line in the figure. We find that only $\approx 8\%$ of the galaxies host a CRD with a mass fraction larger than $1\%$. The sample of galaxies that host a significant CRD, i.e., $M_{\rm CRD}/M_{\rm Disk}\geq 0.01$ consists of 20 galaxies. In this study, we decided to include in our final sample six boundary cases with $0.0095\leq M_{\rm CRD}/M_{\rm Disk}<0.01$. The mass distribution of the final sample is shown in Figure 1 with orange histograms.
In Table 1 we list the internal properties of the galaxy sample with identified CRDs. In the left-most column, the label given to each identified CRD is shown. The corresponding Subfind ID at $z=0$ is given in the second column. In Figure 3 the circularity parameter distribution for eight galaxies is shown. The top four examples in the figure show galaxies that do not host a CRD, while the bottom four examples show galaxies hosting a CRD, according to our definition. The red and light blue shaded regions represent the star particles belonging to the CRD and stellar disk, respectively.
Figure 4 presents eight examples of galaxies hosting a significant CRD. The figure shows their total stellar surface mass density both in face-on and edge-on projections. We show the most extensive CRDs in terms of their $R_{95}^{\rm CRD}$. We note that galactic disks present a morphological diversity, with some exhibiting well-defined bars (e.g., CR-5) and spiral structures (e.g., CR-10 and CR-18).
Figure 5 shows the spatial distribution of the counterrotating stellar particles considered as members of the CRD in the same order as in Fig. 4. By comparing the two figures, we observed that in some galaxies (e.g., CR-5 and CR-3), the CRDs extend to radial distances comparable to those of the overall stellar disk. In contrast, other cases (e.g., CR-25) exhibit much more compact CRDs. In addition, some cases such as CR-0 and CR-5 display an interesting torus-like morphology. We explore the origin of their structure in the following sections. The full set of 26 galaxies analyzed is shown in Fig. 12 and 13.
From this analysis, and based on our criteria, it follows that approximately 10% (including the boundary cases) of the galaxies host a significant stellar CRD. Our results indicate that CRDs are relatively infrequent. This is consistent with previous observational studies, which report CRDs, specially stellar, as rare systems found in only $\approx 1\%$ of galaxies found in Mapping Nearby Galaxies at APO (MaNGA) (Bao et al. 2022, Gasymov et al. 2025).
@@ 4 Characterization of CRDs
Figure 6 shows the relationship between the CRD mass fraction, $M_{\rm CRD}/M_{\rm Disk}$, and different measurements of the disk extension and mass distribution. In particular, we focus on $R_{50}^{\rm CRD}/R_{\rm opt}$ (top), $R_{95}^{\rm CRD}/R_{\rm opt}$ (middle), and $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}$ (bottom). Data points are color coded according to the total stellar mass of the disk.
@@
Figure 6: Coplanar counterrotating disk (CRD) mass fraction, $M_{\rm CRD}/M_{\rm Disk}$, as a function of different measurements of the counterrotating stellar disk extension and mass distribution for the sample of 26 significant CRD galaxies. Top and middle panels: Radius containing 50% and 95% of the counterrotating stellar disk mass normalized by the optical radius of the galaxy, respectively. Bottom panel: Stellar concentration of the coplanar counterrotating disk component. All data points are color coded by the total stellar disk mass. We observed that while most CRDs are compact, with the majority having $R_{50}^{\rm CRD}/R_{\rm opt}<0.25$, there are a few extended cases. Overall, we find that our sample of counter rotating galaxies show a variety of structural properties.
@@
The top panel of Fig. 6 shows that most CRDs are relatively compact in nature, with approximately 88% having a half-mass radius ($R_{50}^{\rm CRD}$) smaller than a quarter of their optical radius ($R_{\rm opt}$). However, the middle panel of the figure reveals that around 89% of CRDs have $R_{95}^{\rm CRD}/R_{\rm opt}$ values greater than 0.25, and approximately 19% exceed half the optical radii. This indicates that while most CRDs are compact in terms of their half-mass radii, many have a reasonably extended radial distribution. The bottom panel of Fig. 6 shows a measure of the CRDs concentration, which is defined as the ratio of $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}$. The CRDs with $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}\leq 2$ contain an approximate significant portion of their mass in a thin ring-like structure. We identified three of these torus-like CRDs, and they are CR-0, CR-5, and CR-22. Indeed, this can be appreciated on Fig. 5. The CRDs with higher concentration values, such as CR-3 and CR-10, exhibit less compact profiles. The wide spread in the CRD concentration highlights the structural diversity within the sample. Overall, we find that our sample of counter rotating galaxies shows a variety of structural properties. Three CRDs show a ring-like structure, some display an extended CRD, and others have a more compact nature with values of $R_{95}^{\rm CRD}$ below half of the disk’s $R_{\rm opt}$.
@@
Next we characterize the origin of the stellar particles that conform each CRD. We recall that stellar particles formed from the gas distribution bound to the potential well of the main host galaxy are defined as in situ. This includes gas that may be recently stripped from closely interacting satellite galaxies. In Fig. 7, we show the fractions of in situ and ex situ stellar particles in each significant CRD. The light blue (bottom) fraction of each bar indicates the contribution from in situ stellar particles, while the other colored fractions indicate the contribution to the ex situ stellar particles from different satellite galaxies. Interestingly, a majority of the CRDs present a dominant in situ contribution, with 19 out of 26 (73%) cases displaying fractions above 80%. The remaining seven CRDs present a more significant ex situ contribution, i.e., $20\%$. We found only one CRD dominated by the ex situ component (CR-6). In general, CRDs with significant ex situ contributions have only one significant progenitor that contributes over 90% of the ex situ particles, with CR-2 being the only exception. We explore these cases further in Section 5.
@@
Figure 7: Fraction of the in situ and ex situ stellar particles in the coplanar counterrotating disk component for the sample of 26 significant CRD galaxies. Light blue bars represent the fraction of in situ stellar particles in the CRD, while the remaining colors represent the contribution to the ex situ fraction in the CRD from different satellite galaxies. Each color refers to a different satellite. The dashed red line indicates an in situ contribution equal to 80%.
@@
In general, we find that CRDs have a very significant in situ contribution and, in most cases, present a compact distribution. We note, however, that some show extended configurations. As we show in Section 5, even though most CRD stellar particles are born within the host potential well, the corresponding star-formation bursts are typically associated with significant interaction with massive nearby satellites.
@@
In Fig. 8 we show the distribution of CRDs in a space defined by the fraction of ex situ material and the extension of the CRD with respect to the overall stellar disk, $R_{95}^{\rm CRD}/R_{\rm opt}$. Each symbol represents a single CRD. The top histogram in the figure highlights that a significant fraction of CRDs ($\approx 80\%$) fall into a compact configuration, with values of $R_{95}^{\rm CRD}/R_{\rm opt}<0.5$. Indeed, the median value of the distribution is $R_{95}^{\rm CRD}/R_{\rm opt}\sim 0.37$. However, we note the presence of the five interesting cases with very extended in situ disks, i.e., $R_{95}^{\rm CRD}/R_{\rm opt}>0.6$ (see also Fig. 6). As previously discussed, the histogram in the right panel of the figure shows that most CRDs ($\approx 73\%$) have an ex situ fraction of less than $0.2$. Indeed, the median value of the ex situ mass fraction is approximately $0.06$.
@@
Based on the space represented by $R_{95}^{\rm CRD}/R_{\rm opt}$ and the ex situ fraction, we classified the CRDs into different categories using a threshold of $R_{95}^{\rm CRD}/R_{\rm opt}=0.5$ and an ex situ fraction of 20%. The different categories are defined as (i) compact in situ, (ii) compact ex situ, and (iii) extended in situ. These categories reflect both their spatial extent and whether the ex situ mass fraction is significant. The classification scheme is highlighted with different colored regions in Fig. 8. Using this scheme, we found that compact in situ CRDs are the most common, with 14 cases, followed by seven compact ex situ and five extended in situ CRDs. Note that no CRDs with a significant ex situ fraction were found to be extended.
@@
This analysis indicates that CRDs are not only rare in TNG50, but they also predominantly exhibit compact spatial distributions. Interestingly, despite being less frequent, it is the extended types that are more likely to be detected observationally, providing context for the $\approx 1\%$ occurrence reported in the MaNGA survey (Bao et al. 2022, Gasymov et al. 2025). Using this CRD classification as a foundation, in the next section we proceed to analyze each type individually, examining their star formation histories (SFHs) and merger activity in order to characterize the nature of their formation.
@@
Figure 8: Classification of the sample of 26 significant CRDs according to their counterrotating stellar disk extension ($R_{95}^{\rm CRD}/R_{\rm opt}$) and to their ex situ fraction of stellar particles. We classified the CRDs into four different types, namely, extended ex situ (yellow quadrant), compact ex situ (red rombus quadrant), compact in situ (blue triangle quadrant) and extended in situ (orange circle quadrant).
@@ 5 Formation channels for CRDs
In this section we explore the different formation channels of the 26 significant CRDs, subdivided according to the three categories defined in Section 4. In our analysis we incorporate information about the stellar ages as well as the properties of infalling satellites.
@@
Figure 9 shows the SFHs of all 26 significant CRDs. Each panel represents a galaxy from our sample. The orange and the light blue lines in the figure represent the SFH of the CR ($\epsilon<-0.7$) and the corotating ($\epsilon>0.7$) disk component, respectively. The vertical dashed lines indicate the median of each distribution. Additionally, the dashed gray line represents the SFH of all stellar particles confined within a cylinder of radius $R=R_{\rm opt}$ and height $|z|\leq 5$ kpc. All SFHs have been normalized to the peak value of the corresponding distribution. It can be clearly seen that with respect to the full disks, CRDs typically have a more bursty history. In most cases, these bursts are also noticeable in the overall disk’s SFH. Additionally, we also noticed that CRDs are significantly older than the corotating counterparts. Indeed, this is the case for $\approx 93\%$ of the identified CRDs. The two exceptions are CR-0 and CR-11. We study these cases further later in this section.
@@
Figure 9 also highlights the more significant interaction events each galaxy has undergone. We focus on interactions associated with satellites that have a total peak mass of at least $5\times 10^{9}$ M and have a satellite to host mass ratio larger than $20\%$ at the corresponding peak mass time. In the figure, the first pericenter passage of such interactions is indicated with arrows, color coded according to the peak mass of each satellite. For CRDs with an ex situ fraction of $\gtrsim 10\%$, we highlight the arrow associated with the most significant ex situ contributor. As expected, in all cases we observed a higher frequency of significant mergers at early times. Nonetheless, significant interactions can take place at any time. Such interactions are commonly associated with star-formation bursts in the overall stellar disk (see also Gargiulo et al. 2017).
5.1 Compact in situ CRDs
@@
Compact CRDs are the most frequent type in our sample. By definition, these systems are not very extended (their values of $R^{\rm CRD}_{95}$ lying under half of the optical radius of the overall disk), meaning the counterrotating stars are largely confined to the spatial limits of the galaxy’s bulge or bar. Within this category, we first considered the CRDs that formed predominantly in situ (ex situ fraction ¡ 20%). There are 14 such cases in our sample. In Fig. 9, the compact in situ CRDs are indicated by blue boxes as well as upward-pointing blue triangle symbols in the top-right corner of each panel.
@@
Figure 9: Star formation histories of all 26 significant CRDs, grouped according to the classification type, namely, compact in situ, compact ex situ, and extended in situ (see Sec.4). The orange and the light blue lines represent the SFH of the counterrotating and the corotating disk component, respectively. The vertical dashed lines indicate the median of each distribution. The gray line represents the SFH of all stellar particles confined within the disk region independent of their circularity parameter. All SFHs were normalized to the peak value of the corresponding distribution. For each CRD, we also label the time of the first pericenter for the most significant interactions experienced by the galaxy, considering only satellites with a peak total mass greater than $5\times 10^{9}M_{\odot}$ and a satellite-to-host mass-ratio greater than 20%. The most significant interaction of the CRD is identified by an arrow outlined in black and white.
@@
Among the compact in situ CRDs, one galaxy, CR-14, stands out as a borderline case. About 18% of its CRD stellar particles are ex situ (accreted). CR-14 lies just below our threshold for the ex situ category, and its counterrotating component can mainly be associated with a significant accretion event. Indeed, Fig. 9 shows that a merger with a satellite with a peak mass on the order of $\approx 8\times 10^{10}$ M induces a star-formation burst that gives rise to most of the corotating component. Furthermore, the assembly information shown in Fig. 7 for the CRD of CR-14 shows that the majority of its few accreted stars came from this single infalling satellite (see Fig. 9, blue highlighted arrow), which contributed stars and gas and induced a burst of star formation.
@@
Many of the compact in situ CRDs show clear signs of interaction-triggered star formation in their histories, even though the stars themselves formed in situ. For example, CR-11, CR-12, CR-14, CR-18, CR-20, and CR-25 all exhibit enhanced star formation in their CRDs that was preceded by an interaction event either at intermediate (e.g., CR-12 and CR-20) or at late times (CR-11). In these galaxies, a passing or merging satellite disturbed the gas in the host, inducing a burst of star formation in the inner galactic region that contributed stellar particles on retrograde orbits. In each of these cases, the CRD’s stellar mass is almost entirely in situ, and the interactions served as a catalyst for star formation rather than directly depositing a large number of stars.
@@
On the other hand, some compact in situ CRDs appear to have formed very early, alongside the initial assembly of the galaxy’s bulge, and without any later significant interactions. Galaxies such as CR-1, CR-8, CR-9, CR-21, CR-22, and CR-24 exhibit strong early bursts of star formation that are occasionally correlated with the high merger rates characteristic of that epoch. During this early phase of galaxy assembly, star formation occurred in a more irregular and turbulent environment, resulting in a pressure-supported stellar population. The more isotropic kinematics of these stellar particles are associated with broad circularity distributions, contributing to both the corotating and counterrotating components of the final stellar disk. Accordingly, the stellar ages of these CRDs trace back to the formation of the primordial inner regions of the galaxies. These structures are therefore old and centrally concentrated.
@@
The two remaining cases have peculiar formation mechanisms. CR-16 formed over a prolonged period during which the corotating component experienced minimal star formation. Until the merger of an infalling satellite, the galaxy lacked a dynamically cold stellar disk. The CRD is thus a remnant of this earlier evolutionary stage when the galaxy was pressure-supported rather than rotation-supported. Finally, CR-13 formed within a short time window following the smooth accretion of misaligned cold gas. The accreted gas triggered a burst of star formation in the central region, giving rise to the CRD.
@@
In summary, the compact in situ CRDs encompass a variety of formation paths. Some (CR-1, CR-8, CR-9, CR-21, CR-22, and CR-24) formed at the earliest times of their host galaxy, essentially as part of the old bulge formation. Others (CR-11, CR-12, CR-18, and CR-20) formed later, with their counterrotating stars born during one or more bursts triggered by satellite interactions (but without significant stellar accretion). A particular case (CR-14) shows a minor contribution from accreted stars in addition to in situ star formation. The CRD of CR-16 is the remnant of a pre-disk stellar population formed prior to the development of a dynamically cold disk, while the CRD of CR-13 originated from the accretion of misaligned gas that fueled central star formation.
@@
From these results, we can broadly categorize the origins of compact in situ CRDs into three scenarios: (i) formation at early times together with the galaxy’s bulge, (ii) interaction-induced star formation without substantial stellar accretion, (iii) peculiar cases associated with smooth misaligned cold gas accretion and pre-stellar disk irregular star formation. The first two mechanisms are the most frequent. In practice, some CRDs may involve a combination of these mechanisms. Overall, however, the common theme is that the compact in situ CRDs were built by star formation within the galaxy and often triggered or influenced by interactions rather than by direct deposition of stars from mergers.
5.2 Compact ex situ CRDs